# IGEV++: Iterative Multi-range Geometry Encoding Volumes for Stereo Matching

### v1

@@ Abstract
Stereo matching is a core component in many computer vision and robotics systems. Despite significant advances over the last decade, handling matching ambiguities in ill-posed regions and large disparities remains an open challenge. In this paper, we propose a new deep network architecture, called IGEV++, for stereo matching. The proposed IGEV++ builds Multi-range Geometry Encoding Volumes (MGEV) that encode coarse-grained geometry information for ill-posed regions and large disparities and fine-grained geometry information for details and small disparities. To construct MGEV, we introduce an adaptive patch matching module that efficiently and effectively computes matching costs for large disparity ranges and/or ill-posed regions. We further propose a selective geometry feature fusion module to adaptively fuse multi-range and multi-granularity geometry features in MGEV. We then index the fused geometry features and input them to ConvGRUs to iteratively update the disparity map. MGEV allows to efficiently handle large disparities and ill-posed regions, such as occlusions and textureless regions, and enjoys rapid convergence during iterations. Our IGEV++ achieves the best performance on the Scene Flow test set across all disparity ranges, up to 768px. Our IGEV++ also achieves state-of-the-art accuracy on the Middlebury, ETH3D, KITTI 2012, and 2015 benchmarks. Specifically, IGEV++ achieves a 3.23% 2-pixel outlier rate (Bad 2.0) on the large disparity benchmark, Middlebury, representing error reductions of 31.9% and 54.8% compared to RAFT-Stereo and GMStereo, respectively. We also present a real-time version of IGEV++ that achieves the best performance among all published real-time methods on the KITTI benchmarks. The code is publicly available at https://github.com/gangweiX/IGEV and https://github.com/gangweiX/IGEV-plusplus.
@@
Large disparity, stereo matching, dense correspondence, cost volume, iterative optimization.
@@ I Introduction
Stereo matching is of substantial interest because it enables the inference of 3D scene geometry from captured images, with applications spanning from 3D reconstruction to robotics and autonomous driving. The key of stereo matching is to find corresponding pixels in the left and right images. The difference between these corresponding pixel locations, known as disparity, can then be used to infer depth and reconstruct the 3D scene. Despite substantial efforts in the field of stereo matching [1, 2, 3, 4, 5, 6, 7, 8], challenges persist in handling occlusions, repetitive structures, textureless or transparent objects. Additionally, effectively managing large disparities in stereo matching remains an open problem.
@@
To overcome the limitations of a fixed disparity range cost volume, STTR [4] adopts a transformer [19] to compute pixel-wise correlation explicitly and densely along epipolar lines. Although STTR [4] has achieved promising results, it is unable to produce predictions for occluded pixels. CSTR [20] attempts to improve STTR’s performance with a context-enhanced transformer architecture, but it still suffers from the same limitations as STTR. GMStereo [5] uses a self-attention layer to propagate predictions from non-occluded regions to occluded regions. Although the transformer-based methods can overcome the limitations of a pre-defined disparity range, they lack an explicit cost volume. As a result, they are unable to integrate multi-scale cost volumes via 3D convolutions and thus have difficulties in resolving ambiguities in ill-posed regions (Figure 4810).
Recently, iterative optimization-based methods [3, 15, 21, 16, 22] have exhibited attractive performance on high-resolution datasets with large disparities. Different from the filtering-based methods, iterative methods avoid the computationally expensive cost aggregation operations and progressively update the disparity map by repeatedly fetching cost information from the all-pairs 4D correlation volume. For instance, RAFT-Stereo [3] calculates all-pairs correlations (APC) between all pixels of the left and right images on the same epipolar lines, and then exploits multi-level Convolutional Gated Recurrent Units (ConvGRUs) [23] to recurrently update the disparity map using local costs retrieved from the APC. Due to the all-pairs 4D correlation volume, RAFT-Stereo can predict large disparities. However, without cost aggregation the original cost volume lacks non-local geometry and context information. As a result, existing iterative methods have difficulties tackling ambiguities in ill-posed regions, such as occlusions, textureless regions and repetitive structures. Although ConvGRUs can improve the predicted disparity map by incorporating context and geometry information from context features and hidden layers, the limitation in the original cost volume greatly limits the effectiveness of each iteration, resulting in the need for a large amount of ConvGRUs iterations to achieve satisfactory performance (Figure 5).
@@
To combine the complementary advantages of filtering-based and iterative optimization-based methods, we propose Iterative Multi-range Geometry Encoding Volumes (IGEV++) which addresses ambiguities in ill-posed regions by aggregating the cost volume using an extremely lightweight 3D regularization network before iterative ConvGRUs optimization. To further address the limitation of the filtering-based method in efficiently handling large disparities, our IGEV++ adopts the novel Multi-range Geometry Encoding Volumes (MGEV), which encode coarse-grained geometry information for ill-posed regions and large disparities, and fine-grained geometry information for details and small disparities (see Figure 3). Our MGEV is inspired by a key observation that objects with small disparities are distant and occupy fewer pixels, while objects with large disparities are nearby and occupy more pixels. We further propose an adaptive patch matching method to enable effective and efficient MGEV construction and a selective geometry information fusion module to efficiently integrate multi-range and multi-granularity information at each iteration.
Our IGEV++ consistently outperforms existing methods [11, 3, 16, 5] by a large margin across all disparity ranges (see Figure 1 and Table I). Specifically, as the disparity range increases, existing methods exhibit a significant drop in accuracy. In contrast, our method remains robust for large disparity ranges. Our method also demonstrates outstanding ability in handling extensive ill-posed regions, achieving the best performance on reflective regions in the KITTI 2012 benchmark. Additionally, the proposed MGEV provides more comprehensive yet concise information for ConvGRUs to update, enabling our IGEV++ to converge faster. For example, our IGEV++ achieves a lower EPE (i.e., 0.79) with only 4 iterations (see Table IV), compared to DLNR’s 32 iterations (i.e., EPE of 0.81 for inference).
To fully exert the advantages of the proposed geometry encoding volume, we also introduce a real-time version of IGEV++, named RT-IGEV++, to offer an attractive solution for time-constrained applications. Our RT-IGEV++ achieves real-time speed and the best accuracy among all published real-time methods [17, 24, 25, 26, 27] on the KITTI benchmarks [12, 13].
@@
  • We propose a real-time version of IGEV++ that can achieve real-time inference and the best performance among all published real-time methods.
@@ II-A Cost Volume Filtering-based Methods
Cost Volume Filtering-based Methods typically consist of four steps, i.e. feature extraction, cost volume construction, cost aggregation and disparity regression. To improve the representative ability of a cost volume, most existing learning-based stereo methods [1, 28, 29, 30, 31, 32, 27, 33, 34, 35, 36, 37, 38, 39] construct a cost volume using powerful CNN features. However, the cost volume could still suffer from the ambiguity problem in occluded regions, large textureless/reflective regions and repetitive structures. The 3D convolutional networks have exhibited great potential in regularizing or filtering the cost volume, which can propagate reliable sparse matches to ambiguous and noisy regions. GCNet [40] firstly uses 3D encoder-decoder architecture to regularize a 4D concatenation volume. PSMNet [1] proposes a stacked hourglass 3D CNN in conjunction with intermediate supervision to regularize the concatenation volume. GwcNet [2] and ACVNet [9, 27] propose the group-wise correlation volume and the attention concatenation volume, respectively, to improve the expressiveness of the cost volume and in turn improve the performance in ambiguous regions. PCWNet [11] propose pyramid combination and warping cost volume to improve cross-domain generalization [41, 42, 43, 44, 45, 46]. GANet [47] designs a semi-global aggregation layer and a locally guided aggregation layer to further improve the accuracy. However, the high computational and memory costs of 3D CNNs often prevent these models from being applied to high-resolution cost volumes. To improve memory efficiency, several cascade methods [10, 48, 27] have been proposed. CFNet [10] and CasStereo [48] build a cost volume pyramid in a coarse-to-fine manner to progressively narrow down the predicted disparity range. Despite their impressive performance, the coarse-to-fine methods inevitably involve accumulated errors at coarse resolutions. In contrast, our IGEV++ can iteratively update disparity at a high resolution through ConvGRU.
@@ II-B Iterative Optimization-based Methods
Recently, several iterative optimization-based methods [49, 3, 50, 15, 51, 16, 52, 53] have been proposed and have achieved impressive performance in dense correspondence tasks. RAFT [49] pioneeringly introduces a deep iterative architecture for optical flow estimation that uses ConvGRU to iteratively update the flow field with high accuracy and robustness. Subsequently, RAFT-Stereo [3] introduces multi-level ConvGRUs to iteratively update the disparity field using local cost values retrieved from all-pairs correlations. CREStereo [15] designs a hierarchical network with recurrent refinement to update the disparity field in a coarse-to-fine manner. To mitigate the loss of detail information during the iterations, DLNR [16] designs the Decouple LSTM module to decouple the hidden state from the update operator. However, these iterative methods mainly rely on the all-pairs correlations (APC) which lack non-local information and thus face challenges in resolving local ambiguities in ill-posed regions. Our IGEV++ also adopts ConvGRUs as RAFT-Stereo [3] to iteratively update the disparity field. But different from RAFT-Stereo [3], we construct MGEV to encode non-local geometry and context information which can significantly improve the effectiveness of each ConvGRUs iteration. In addition, we provide an accurate initial disparity map for the ConvGRUs updater to start at a negligible cost, yielding a much faster convergence than RAFT-Stereo [3].
@@ II-C Transformer-based Stereo Methods
With the rise of attention mechanisms, transformer-based deep stereo methods [4, 20, 54, 5] have become another line of stereo matching research. A representative method is STTR [4], which revisits the stereo matching problem from a sequence-to-sequence correspondence perspective with transformer to replace the fixed-range cost volume. Following this, CSTR [20] introduces a context-enhanced transformer architecture to improve STTR’s performance. GMStereo [5] leverages the transformer to integrate knowledge between the left and right images via cross-view interactions. Recently, Croco-Stereo [6] achieves impressive performance by scaling up vision transformer-based cross-completion architectures with large amounts of data. While transformer-based stereo methods excel in modeling long-range dependencies and can handle large disparities to some extent, they encounter difficulties in addressing local ambiguities in ill-posed regions (such as occlusions and reflective regions) due to the absence of an explicit cost volume.
@@ III-A Multi-scale Feature Extraction
Feature extraction contains two parts: 1) a feature network that extracts multi-scale features for cost volume construction and cost aggregation guidance, and 2) a context network that extracts multi-scale context features for ConvGRUs hidden state initialization and updating.
Feature Network. Given the left and the right images $\mathbf{I}_{l(r)}\in\mathbb{R}^{{3}\times{H}\times{W}}$, we first apply the MobileNetV2 pretrained on ImageNet [55] to scale $\mathbf{I}_{l(r)}$ down to 1/32 of the original size, and then use upsampling blocks with skip-connections to recover them up to 1/4 scale, resulting in multi-scale features $\mathbf{f}_{l,i}(\mathbf{f}_{r,i})\in\mathbb{R}^{{C_{i}}\times{\frac{H}{i}}\times{\frac{W}{i}}}$} ($i$=4, 8, 16, 32 and $C_{i}$ for feature channels). The $\mathbf{f}_{l,4}$ and $\mathbf{f}_{r,4}$ are used to construct the multi-range cost volumes. And the $\mathbf{f}_{l,i}$ ($i$=4, 8, 16, 32) are also used as context guidance for 3D regularization network.
Context Network. Following RAFT-Stereo [3], the context network consists of a series of residual blocks [56] and downsampling layers, producing multi-scale context features at 1/4, 1/8 and 1/16 of the input image resolution with 128 channels. The multi-scale context features are used to initialize the hidden state of the ConvGRUs and are also inserted into the ConvGRUs at each iteration.
@@ III-B Multi-range Geometry Encoding Volumes Construction
Given the left feature map $\mathbf{f}_{l,4}$ and right feature map $\mathbf{f}_{r,4}$ extracted from $\mathbf{I}_{l}$ and $\mathbf{I}_{r}$, We build group-wise correlation volumes respectively for small disparity range $D^{s}$ (<192px), medium disparity range $D^{m}$ (<384px), and large disparity range $D^{l}$ (<768px). The group-wise correlation volume [2] refers to splitting features $\mathbf{f}_{l,4}$ ($\mathbf{f}_{r,4}$) into $N_{g}$ ($N_{g}$=8) groups along the channel dimension and computes correlation maps group by group. For small disparity ranges, the correlation volume $\mathbf{C}^{s}$ is constructed by,
@@
$\mathbf{C}^{l}(g,d^{l},x,y)\!=\!\frac{1}{N_{c}/N_{g}}\sum\limits_{i=0}^{3}\omega_{i}\langle\mathbf{f}_{l,4}^{g}(x,y),\mathbf{f}_{r,4}^{g}(x-(d^{l}+i),y)\rangle,$ (2)
where $\omega_{i}$ represents the weight of a matching cost in the patch and is learned adaptively during the training process, and the candidate disparity range $d^{l}\!\in\!\{0,4,8,\dots,D^{l}-4\}$.

### v2

@@ Abstract
Stereo matching is a core component in many computer vision and robotics systems. Despite significant advances over the last decade, handling matching ambiguities in ill-posed regions and large disparities remains an open challenge. In this paper, we propose a new deep network architecture, called IGEV++, for stereo matching. The proposed IGEV++ constructs Multi-range Geometry Encoding Volumes (MGEV), which encode coarse-grained geometry information for ill-posed regions and large disparities, while preserving fine-grained geometry information for details and small disparities. To construct MGEV, we introduce an adaptive patch matching module that efficiently and effectively computes matching costs for large disparity ranges and/or ill-posed regions. We further propose a selective geometry feature fusion module to adaptively fuse multi-range and multi-granularity geometry features in MGEV. Then, we input the fused geometry features into ConvGRUs to iteratively update the disparity map. MGEV allows to efficiently handle large disparities and ill-posed regions, such as occlusions and textureless regions, and enjoys rapid convergence during iterations. Our IGEV++ achieves the best performance on the Scene Flow test set across all disparity ranges, up to 768px. Our IGEV++ also achieves state-of-the-art accuracy on the Middlebury, ETH3D, KITTI 2012, and 2015 benchmarks. Specifically, IGEV++ achieves a 3.23% 2-pixel outlier rate (Bad 2.0) on the large disparity benchmark, Middlebury, representing error reductions of 31.9% and 54.8% compared to RAFT-Stereo and GMStereo, respectively. We also present a real-time version of IGEV++ that achieves the best performance among all published real-time methods on the KITTI benchmarks. The code is publicly available at https://github.com/gangweix/IGEV and https://github.com/gangweix/IGEV-plusplus.
@@
Large disparity, stereo matching, multi-range, dense correspondence, cost volume, iterative optimization.
@@ I Introduction
Stereo matching is of substantial interest because it enables the inference of 3D scene geometry from captured images, with applications spanning from 3D reconstruction to robotics and autonomous driving. The key to stereo matching is to find corresponding pixels in the left and right images. The difference between these corresponding pixel locations, known as disparity, can then be used to infer depth and reconstruct the 3D scene. Despite substantial efforts in the field of stereo matching [1, 2, 3, 4, 5, 6, 7, 8], challenges persist in handling occlusions, repetitive structures, textureless or transparent objects. Additionally, effectively managing large disparities in stereo matching remains an open problem.
@@
To overcome the limitations of a fixed disparity range cost volume, STTR [4] adopts a transformer [19] to compute pixel-wise correlation explicitly and densely along epipolar lines. Although STTR [4] has achieved promising results, it is unable to produce predictions for occluded pixels. CSTR [20] attempts to improve STTR’s performance with a context-enhanced transformer architecture, but it still suffers from the same limitations as STTR. GMStereo [5] uses a self-attention layer to propagate predictions from non-occluded regions to occluded regions. Although the transformer-based methods can overcome the limitations of a pre-defined disparity range, they lack an explicit cost volume. As a result, they are unable to integrate multi-scale cost volumes via 3D convolutions and thus have difficulties in resolving ambiguities in ill-posed regions (Figures 4912).
Recently, iterative optimization-based methods [3, 15, 21, 16, 22] have exhibited attractive performance on high-resolution datasets with large disparities. Different from the filtering-based methods, iterative methods avoid the computationally expensive cost aggregation operations and progressively update the disparity map by repeatedly fetching cost information from the all-pairs 4D correlation volume. For instance, RAFT-Stereo [3] calculates all-pairs correlations (APC) between all pixels of the left and right images on the same epipolar lines, and then exploits multi-level Convolutional Gated Recurrent Units (ConvGRUs) [23] to recurrently update the disparity map using local costs retrieved from the APC. Due to the all-pairs 4D correlation volume, RAFT-Stereo can predict large disparities. However, without cost aggregation the original cost volume lacks non-local geometry and context information. As a result, existing iterative methods have difficulties tackling ambiguities in ill-posed regions, such as occlusions, textureless regions, and repetitive structures. Although ConvGRUs can improve the predicted disparity map by incorporating context and geometry information from context features and hidden layers, the limitation in the original cost volume greatly limits the effectiveness of each iteration, resulting in the need for a large amount of ConvGRUs iterations to achieve satisfactory performance (Figure 5).
@@
To combine the complementary advantages of filtering-based and iterative optimization-based methods, we propose Iterative Geometry Encoding Volume (IGEV) and Iterative Multi-range Geometry Encoding Volumes (IGEV++), which address ambiguities in ill-posed regions by aggregating the cost volume using an extremely lightweight 3D regularization network before iterative ConvGRUs optimization. To further address the limitation of the filtering-based method and efficiently handle large disparities, our IGEV++ adopts the novel Multi-range Geometry Encoding Volumes (MGEV), which is computed effectively and efficiently via the multi-granularity matching cost calculation approach. Specifically, we compute coarse-grained matching costs (point-to-patch) for the GEV with a large disparity range, and fine-grained matching costs (point-to-point) for the GEV with a small disparity range, as illustrated in Figure 3. MGEV encodes coarse-grained geometry information for ill-posed regions and large disparities, while preserving fine-grained geometry information for detailed structures and small disparities. Our MGEV is inspired by a key observation that objects with small disparities are distant and occupy fewer pixels, while objects with large disparities are nearby and occupy more pixels. We further propose an adaptive patch matching method to enable effective and efficient MGEV construction and a selective geometry feature fusion module to efficiently integrate multi-range and multi-granularity information at each iteration.
Our IGEV++ consistently outperforms existing methods [11, 3, 16, 5] by a large margin across all disparity ranges (see Figure 1 and Table I). Specifically, as the disparity range increases, existing methods exhibit a significant drop in accuracy. In contrast, our method remains robust for large disparity ranges. Our method also demonstrates outstanding ability in handling extensive ill-posed regions, achieving the best performance on reflective regions in the KITTI 2012 benchmark. Additionally, the proposed MGEV provides more comprehensive yet concise information for ConvGRUs to update, enabling our IGEV++ to converge faster. For example, our IGEV++ achieves a lower EPE (i.e., 0.79) with only 4 iterations (see Table V), compared to DLNR’s 32 iterations (i.e., EPE of 0.81 for inference).
To fully exert the advantages of the proposed geometry encoding volume, we also introduce a real-time version of IGEV++, named RT-IGEV, to offer an attractive solution for time-constrained applications. Our RT-IGEV achieves real-time speed and the best accuracy among all published real-time methods [17, 24, 25, 26, 27] on the KITTI benchmarks [12, 13].
@@
  • We propose RT-IGEV, a real-time version of IGEV++, which achieves real-time inference while delivering the best performance among all published real-time methods.
@@ II-A Cost Volume Filtering-based Methods
Cost Volume Filtering-based Methods typically consist of four steps, i.e. feature extraction, cost volume construction, cost aggregation and disparity regression. To improve the representative ability of a cost volume, most existing learning-based stereo methods [1, 28, 29, 30, 31, 32, 27, 33, 34, 35, 36, 37, 38, 39] construct a cost volume using powerful CNN features. However, the cost volume could still suffer from the ambiguity problem in occluded regions, large textureless/reflective regions and repetitive structures. The 3D convolutional networks have exhibited great potential in regularizing or filtering the cost volume, which can propagate reliable sparse matches to ambiguous and noisy regions. GCNet [40] firstly uses 3D encoder-decoder architecture to regularize a 4D concatenation volume. PSMNet [1] proposes a stacked hourglass 3D CNN in conjunction with intermediate supervision to regularize the concatenation volume. GwcNet [2] and ACVNet [9, 27] propose the group-wise correlation volume and the attention concatenation volume, respectively, to improve the expressiveness of the cost volume and in turn improve the performance in ambiguous regions. PCWNet [11] propose pyramid combination and warping cost volume to improve cross-domain generalization [41, 42, 43, 44, 45, 46]. GANet [47] designs a semi-global aggregation layer and a locally guided aggregation layer to further improve the accuracy. However, the high computational and memory costs of 3D CNNs often prevent these models from being applied to high-resolution cost volumes. To improve memory efficiency, several cascade methods [10, 48, 27] have been proposed. CFNet [10] and CasStereo [48] build a cost volume pyramid in a coarse-to-fine manner to progressively narrow down the predicted disparity range. Despite their impressive performance, the coarse-to-fine methods inevitably involve accumulated errors at coarse resolutions. In contrast, our IGEV++ can iteratively update disparity map at a high resolution through ConvGRUs.
@@ II-B Iterative Optimization-based Methods
Recently, several iterative optimization-based methods [49, 3, 50, 15, 51, 16, 52, 53] have been proposed and have achieved impressive performance in dense correspondence tasks. RAFT [49] pioneeringly introduces a deep iterative architecture for optical flow estimation that uses ConvGRU to iteratively update the flow field with high accuracy and robustness. Subsequently, RAFT-Stereo [3] introduces multi-level ConvGRUs to iteratively update the disparity field using local cost values retrieved from all-pairs correlations. CREStereo [15] designs a hierarchical network with recurrent refinement to update the disparity field in a coarse-to-fine manner. To mitigate the loss of detail information during the iterations, DLNR [16] designs the Decouple LSTM module to decouple the hidden state from the update operator. However, these iterative methods mainly rely on the all-pairs correlations (APC) which lack non-local information and thus face challenges in resolving local ambiguities in ill-posed regions. Our IGEV++ also adopts ConvGRUs as RAFT-Stereo [3] to iteratively update the disparity map. But different from RAFT-Stereo [3], we construct MGEV to encode non-local geometry and context information which can significantly improve the effectiveness of each iteration. In addition, we provide an accurate initial disparity map for the ConvGRUs updater to start at a negligible cost, yielding a much faster convergence than RAFT-Stereo [3].
@@ II-C Transformer-based Stereo Methods
With the rise of attention mechanisms, transformer-based deep stereo methods [4, 20, 54, 5] have become another line of stereo matching research. A representative method is STTR [4], which revisits the stereo matching problem from a sequence-to-sequence correspondence perspective with transformer to replace the fixed-range cost volume. Following this, CSTR [20] introduces a context-enhanced transformer architecture to improve STTR’s performance. GMStereo [5] leverages the transformer to integrate knowledge between the left and right images via cross-view interactions. Recently, Croco-Stereo [6] has achieved impressive performance by scaling up vision transformer-based cross-completion architectures with large amounts of data. While transformer-based stereo methods excel in modeling long-range dependencies and can handle large disparities to some extent, they encounter difficulties in addressing local ambiguities in ill-posed regions (such as occlusions and reflective regions) due to the absence of an explicit cost volume.
@@ III-A Multi-scale Feature Extraction
Feature extraction contains two parts: 1) a feature network that extracts multi-scale features for cost volume construction and cost aggregation guidance, and 2) a context network that extracts multi-scale context features for ConvGRUs hidden state initialization and update.
Feature Network. Given the left and the right images $\mathbf{I}_{l(r)}\in\mathbb{R}^{{3}\times{H}\times{W}}$, we first apply the MobileNetV2 pretrained on ImageNet [55] to scale $\mathbf{I}_{l(r)}$ down to 1/32 of the original size, and then use upsampling blocks with skip-connections to recover them up to 1/4 scale, resulting in multi-scale features $\mathbf{f}_{l,i}(\mathbf{f}_{r,i})\in\mathbb{R}^{{C_{i}}\times{\frac{H}{i}}\times{\frac{W}{i}}}$} ($i$=4, 8, 16, 32 and $C_{i}$ for feature channels). The $\mathbf{f}_{l,4}$ and $\mathbf{f}_{r,4}$ are used to construct the multi-range cost volumes. And the $\mathbf{f}_{l,i}$ ($i$=4, 8, 16, 32) are also used as context guidance for cost volume filtering.
Context Network. Following RAFT-Stereo [3], the context network consists of a series of residual blocks [56] and downsampling layers, producing multi-scale context features at 1/4, 1/8 and 1/16 of the input image resolution with 128 channels. The multi-scale context features are used to initialize the hidden state of the ConvGRUs and are also inserted into the ConvGRUs to update the disparity map at each iteration.
@@ III-B Multi-range Geometry Encoding Volumes Construction
Given the left feature map $\mathbf{f}_{l,4}$ and right feature map $\mathbf{f}_{r,4}$ extracted from $\mathbf{I}_{l}$ and $\mathbf{I}_{r}$, we construct group-wise correlation volumes respectively for small disparity range $D^{s}$ (<192px), medium disparity range $D^{m}$ (<384px), and large disparity range $D^{l}$ (<768px). The group-wise correlation volume [2] refers to splitting features $\mathbf{f}_{l,4}$ ($\mathbf{f}_{r,4}$) into $N_{g}$ ($N_{g}$=8) groups along the channel dimension and computes correlation maps group by group. For small disparity ranges, the correlation volume $\mathbf{C}^{s}$ is constructed by,
@@
$\mathbf{C}^{l}(g,d^{l},x,y)\!=$ $\;\!\frac{1}{N_{c}/N_{g}}\langle\mathbf{f}_{l,4}^{g}(x,y),\mathbf{p}^{g}(x-d^{l},y)\rangle,$ (2)
$\mathbf{p}^{g}(x-d^{l},y)=$ $\;\sum\limits_{i=0}^{3}\omega_{i}\mathbf{f}_{r,4}^{g}(x-(d^{l}+i),y)$
where the candidate disparity values $d^{l}\in\mathcal{D}^{l}$, $\mathcal{D}^{l}=\{0,4,8,\dots,D^{l}/4-4\}$, consisting of 48 disparity candidates (768/4/4) for cost aggregation. $\mathbf{C}^{l}$ is constructed by point-to-patch matching (coarse-grained matching) between the left and right images (Figure 3). $\omega_{i}$ represents the weight of the matching cost in a patch of the right image and is learned adaptively during the training process. To efficiently handle large disparities, we propose adaptive patch matching to represent a large disparity range with fewer disparity candidates, significantly reducing the computational cost of cost aggregation and lowering the difficulty of disparity regression.